# More Modality, More AI: Exploring Design Opportunities of AI-Based Multi-modal Remote Monitoring Technologies for Early Detection of Mental Health Sequelae in Youth Concussion Patients

### v1

@@ 6.2.2. The Sleep and Physical Activities Module
In this module (Fig.2 B) concussion clinicians can hover over the question mark icon next to the module title to view its data source, which comes from wearable devices. Below the title, there are options to view data from the past week or the most recent day, with the default display set to one week of data. The module includes three selectable tabs: Sleep Data, Physical Activities, and Recommendation Compliance. The displayed information updates dynamically based on the selected tab. In the Sleep Data tab, clinicians can access heart rate, nap duration, sleep and wake-up times, and total sleep hours. They can select one or multiple metrics based on their specific needs. Different data types are displayed in distinct colors to help clinicians visually distinguish between them more easily. When hovering over a data point in the chart, the exact value for that day is displayed, providing detailed insights into the patient’s sleep patterns.
6.2.3. The Self-Report Symptoms Module
On the right side of the module (Fig.2 C) title, concussion clinicians can hover over an icon to view the data source, which comes from CAs. The module provides two options for selecting a time range: the past week and the last 24 hours. Below the time selector, the interface is divided into two sections. On the left side, there are two tabs: one displaying VA-patient conversation history and the other showing VA-parent conversation history. Clinicians can expand the conversation panel by clicking the button in the top right corner to focus on historical dialogues. On the right side, the Symptoms Reported section visualizes symptom-related information. If the patient or parent has mentioned any symptoms, a colored indicator dot appears next to the corresponding symptom category, with the color reflecting severity—red for severe symptoms, and orange for moderate symptoms. If the patient reports no symptoms, the corresponding dot turns green. To the right of each category, the symptom duration is displayed, indicating how long the reported issue has persisted. Clinicians can also click AI Summary to view a detailed breakdown of the category, providing additional context to support their assessment. At the bottom of the Symptoms Reported section, we provide call buttons for contacting the patient or parent, which ensures that clinicians can quickly intervene in case of an emergency.
6.2.4. The Questionnaires Result Module
@@
In this module (Fig. 2 D), the title includes an option to view data source information, which comes from patient-administered questionnaires at home. These questionnaires include GAD-7 for anxiety, PHQ-9 for depression, and PCSS for concussion severity. The middle section of the chart presents the scores for each questionnaire along with corresponding explanations to help clinicians interpret the results. The module also displays a comparison with previous scores, providing clinicians with a sense of the patient’s recovery trend. Additionally, clinicians can click ”Show Details” to view a breakdown of individual section scores within each questionnaire for a more in-depth assessment.
6.2.5. The AI Risk Score Prediction Module
The final module (Fig. 2 F) presents the AI-generated prediction of a patient’s risk of developing mental health sequelae. Concussion clinicians can hover over the question mark icon next to the title to view information about the source of the risk score. Based on the current predictive capabilities of the AI model, this module offers four different time range options, with the default set to risk predictions within four weeks. Below the time selector, clinicians can view the risk score chart for the selected period and hover over specific time points to examine the risk score at each point. At the bottom of the module, we display the features contributing to the risk score, along with the influence of each contributing factor, represented as a percentage. Visually, if a feature has a high contribution to the risk score, we provide a visual alert. The dot next to the feature title changes color based on the severity of the symptoms associated with that feature, with red indicating severe symptoms and orange indicating moderate severity. This severity level is also reflected in the percentage value on the right, using corresponding colors to enhance clarity. Additionally, the visual representation of each contributing module is color-matched to its corresponding section in the UI. This alignment helps clinicians intuitively understand the connections between different modules and their impact on the AI risk score.
@@ 7. Discussion
In this work, we identified multiple challenges that concussion clinicians face when treating youth concussion patients with mental health sequelae, including the difficulties of collecting patients’ information outside the clinics, communicating with patients and their families, as well as making timely clinical decision-making about concussion treatment based on mental health symptoms. Further, we proposed multiple AI-RPM technologies accompanied by AI prediction models to support concussion clinicians’ clinical workflow and iteratively refined a clinician-oriented EMR dashboard user interface to consolidate detailed feedback on information representations, interactions, customizability, and multiple responsible AI considerations. In this section, we delve into how the aforementioned challenges arise and explore how they can potentially be addressed through AI-RPM technologies. Additionally, we discuss the role of AI-RPM systems in supporting clinicians’ decision-making processes. Finally, we talk about concerns regarding responsibility in the potential use of AI-RPM systems and identify design opportunities to tackle these challenges.
7.1. Remote Monitoring of Patient Mental Health-Related Nonclinical Data
7.1.1. Concussion Clinicians’ Need of Monitoring Patients’ Mental Health Symptoms
A unique challenge faced by concussion clinicians, which was identified in our work, is that clinicians find it critical but difficult to assess the severity of patients’ mental health symptoms when they are outside the clinic. For concussion patients, there is a significant overlap between concussion symptoms and mental health symptoms, and these two illnesses could influence each other’s recovery progress. Although concussion clinicians do not possess expertise in the diagnosis and treatment of mental health issues, their clinical decision-making for concussion treatment is closely correlated with patients’ mental health symptoms. If patients exhibit severe mental health symptoms, concussion clinicians need to adjust the concussion treatment plan and the scheduling for follow-up visits accordingly and refer the patients to mental health specialists for professional intervention.
@@ 7.1.1. Concussion Clinicians’ Need of Monitoring Patients’ Mental Health Symptoms
In clinical scenarios, it is not uncommon to have multiple illnesses intertwined with each other. For instance, a recent work discovered that cancer oncologists need to be constantly aware of patients’ cardiotoxicity during cancer treatment and collaborate with cardiologists to adjust the cancer treatment plan (Wu et al., 2024). Nevertheless, these two scenarios are fundamentally different. Based on our formative study, concussion clinicians do not collaborate with mental health specialists, and neither do they diagnose mental health disorders. Concussion clinicians, in most of their clinical practices, only need to detect whether patients’ mental health issues surpass a certain level of severity threshold. If the mental health issues are severe, concussion clinicians will refer patients to mental health specialists for mental health intervention. Otherwise, concussion clinicians will not do the referral but take the patient’s mental health status as a factor in the design of concussion recovery recommendations.
We believe the integrated design of leveraging AI-RPM technologies to collect patients’ mental health-related data outside the clinic and then developing AI-based predictive models to assess patients’ mental health symptoms could provide actionable insights into concussion clinicians’ current workflow. As suggested by Bennett et al. (2012); Zhang et al. (2024), clinicians expect AI predictive models to provide actionable insights, which is critical to ensure AI models’ usefulness in clinicians’ workflow. Future research could further explore the design of AI-RPM technologies with respect to what other modalities of patients’ data could be beneficial for mental health assessment and further develop AI models to integrate such data, along with contextual information like medical histories and behavioral patterns for more reliable predictions with sufficient explainabilities.
7.1.2. Remote Monitoring of Objective Patient Data
Our study revealed significant challenges in how concussion clinicians assess mental health issues in youth concussion patients. Youth patients often avoid openly discussing mental health concerns due to the stigma of being labeled with mental health disorders. Consequently, clinicians are often limited to collecting patient information through mental health assessment questionnaires or direct observation of patients’ behaviors during clinical visits. However, if patients are unwilling to communicate openly, the validity of questionnaires can be compromised, and observation during clinical visits can be left as the only viable option. However, direction observation is inherently limited from the geographical perspective. Once patients leave the clinic, concussion clinicians can hardly track patients’ mental health status in a timely and effective manner. Even if patients are willing to faithfully and openly discuss their mental health status, patients’ memories could be fragmented and inaccurate, especially for sleeping patterns.
We suggest that RPM technologies, particularly wearable devices like smartwatches, could help address the aforementioned limitations by providing clinicians with real-time access to key physiological metrics. In particular, these physiological metrics are objective measurements that can faithfully reflect patients’ mental health status. Prior research demonstrated the feasibility and advantages of wearable in monitoring objective patient information outside the clinic (Mendel et al., 2024; Frazier et al., 2023; Su et al., 2019; Wu et al., 2024). Wearable devices could monitor objective data such as sleep patterns, physical activity levels, and resting heart rate, which are critical signals for tracking and assessing patients’ mental well-being. For instance, elevated anxiety levels may be inferred through reduced physical activity or disrupted sleep patterns, while a stable resting heart rate could indicate recovery progress. Concussion clinicians could use such data for clinical decision-making, like scheduling timely follow-up visits or initiating referrals to mental health specialists as needed.
An additional benefit provided by wearable devices is that they can monitor patients’ status non-intrusively to be seamlessly integrated into patients’ daily routines. This non-intrusive characteristic minimizes potential cognitive burdens added to the patients, where patients will only need to remember to charge the wearables regularly.
7.1.3. Remote Monitoring of Subjective Patient Data
Clinicians in our study emphasized that the severity and duration of symptoms such as anxiety, depression, or even suicidal ideation are critical factors in assessing whether youth concussion patients are experiencing mental health sequelae. However, these symptoms are more of subjective experiences that can hardly be captured using wearables because the data can be collected by wearables are standardized objective physiological measurements.
To address this gap, we propose integrating advanced generative AI with RPM technology to gather patients’ subjective data through text-based or voice-based conversational assistants (CAs). These CAs are supported by large language models (LLMs) that could engage in free-form natural language conversations with patients. Compared with traditional RPM approaches, natural conversation significantly reduces the requirement of patients’ technical literacy to engage with the devices, enables flexible patient self-reporting, and can be further designed to conduct regular health check-ins through pre-set question lists. Recently, a number of works have demonstrated the promising opportunities of LLM-based CA to support various patient populations in healthcare scenarios (Bartle et al., 2023, 2022; Wang et al., 2023a; Simpson et al., 2020; Ma et al., 2024; Yang et al., 2023), for example, to support older adults with chronic diseases (Yang et al., 2024).
Using AI-driven analytics, key psychological risk factors such as suicidal thoughts could be automatically detected. LLMs could then generate a summary report of the patient’s mental health for the day and transmit it remotely to the clinicians’ EMR system. By combining subjective data from CAs with objective information collected by wearables, clinicians could gain a more comprehensive understanding of the patients’ mental health in their daily lives. Subjective information often provides contextual insights for objective data, which enables clinicians to make more accurate judgments. For example, if wearable data shows that a patient had very low physical activity on a given day, the CA might reveal through its conversation that the patient’s mental health was stable and the low activity was due to a family gathering at home. This additional context would allow the clinician to assess the situation more accurately, avoiding unnecessary concern over what might otherwise appear to be a worrisome pattern in objective metrics. By leveraging the strengths of both data types, our proposed AI-RPM approach could enhance the quality and precision of mental health assessments in youth concussion care.
7.1.4. Patient Privacy Consideration With AI-RPM Technologies
Lastly, when designing such an AI-RPM system, it is essential to underscore protecting patient privacy and ensuring data security. This issue has been widely discussed in many studies and is recognized as a core challenge in AI system design (Gawankar et al., 2024; Bala et al., 2024; Korobenko et al., 2024). However, in our research context, the target population comprises youth patients aged 11–17, whose privacy needs are uniquely sensitive. The privacy of youth patients is not only a concern for the patients themselves but is also critical to their parents, who often play a key role in the medical process. Parents typically expect to have transparency and informed consent regarding how their child’s data is collected and used (Sisk et al., 2020; Haley et al., 2024). Therefore, finding ways to protect patients’ privacy while building and maintaining parental trust in the data collection and usage process is a critical issue that requires careful consideration and balance. Future research could explore methods to effectively safeguard youth health data while simultaneously building trust and participation among parents.
7.2. Multi-Stakeholder Collaboration in the Design of Personalized and Accountable AI-RPM Systems
7.2.1. Collaboration With Clinicians in Supporting Clinical Workflow
Our findings highlight that while clinicians are open to using AI-RPM systems, they are concerned about the potential for information overload. Clinicians are already overwhelmed by the variety of data and tasks within existing EHR systems. Introducing another AI system with additional data could significantly increase their workload, potentially reducing the number of available weekly appointment slots and further straining already limited healthcare resources. In our study, clinicians generally believe they do not have the time to process such detailed information in their daily work. However, clinicians still expressed willingness to check the data collected by the AI-RPM systems before each clinical visit because they believe such data could support them better prepare for patient consultations.
To address this concern, AI-RPM systems need to be designed in collaboration with clinicians to prioritize personalized and context-relevant data presentation. One approach to achieving this is through layered information presentation, where critical data is prioritized and displayed prominently, while secondary information remains accessible but hidden by default. For example, clinicians could be shown a summary of essential metrics, such as daily activity levels, sleep patterns, and key symptom trends, while additional context or raw data could be available on demand (e.g., through interactive elements on the interface). Designers must collaborate closely with clinicians to determine which data points should be prioritized and how they should be presented to support efficient clinical decision-making.
Clinicians in our study also emphasized the importance of defining thresholds for symptom severity, particularly in distinguishing between mild, moderate, and severe symptoms. Some concussion clinicians mentioned that they might not be particularly concerned about mild anxiety, but severe or prolonged symptoms would require their attention. Without clear definitions, AI systems may either underreport or overemphasize symptoms, leading to delayed interventions or unnecessary concerns. Clinicians are the domain experts who can help designers bridge the gap with their knowledge and expertise. By working with clinicians to establish these thresholds, AI-RPM systems can more effectively categorize symptoms, ensuring that clinicians are alerted only when necessary. Such a collaborative approach will not only reduce cognitive overload but also build trust and confidence in the system’s reliability, which will allow clinicians to rely on the AI system as a supportive tool rather than a burden.
Ultimately, the goal of engaging concussion clinicians in the collaborative design process is to enable the system for clinicians to access the most relevant patient information quickly, improve the efficiency of their workflows, and make informed decisions without being overwhelmed by data. By integrating clinicians’ preferences and needs into the system design, AI-RPM systems can enhance their usability and effectiveness, promoting better human-AI collaboration in clinical practice.
7.2.2. Collaboration with Parents in Managing Emergency Situations
In the proposed design section of self-reported symptoms, there is a ”suicidal ideation” alert module which clinicians believe would be of significant benefit to alert clinicians for any unexpected emergency. However, clinicians expressed considerable concerns about this feature. Clinicians questioned the accuracy of the ”suicidal ideation” alerts, not solely in terms of the technical capabilities of the AI system but more importantly regarding the reliability of the self-reported data. In our research context, clinicians pointed out that at the developmental stage, youth patients may often express emotions with exaggeration. For example, youth patients might say things like ”I want to hurt myself” or ”I have suicidal thoughts,” which might not necessarily reflect their actual intentions. Such statements could result in a false alarm when captured by the AI system and misclassified as a ”high-risk alert.” The implications of such false alarms are multifaceted.
First, responding to a high-risk alert would require clinicians to dedicate considerable time and resources to contact the patient and address the situation, even if it turns out to be unwarranted. False alarms could lead to a waste of a clinician’s valuable time and critical medical resources that could be allocated to other patients. Second, if false alarms occur frequently, clinicians could gradually lose trust in the system, which eventually leads to the abandonment of such a system. Another concern expressed by clinicians is the ambiguity of responsibility that AI systems introduce in medical decision-making. Previous studies have highlighted the phenomenon where clinicians, as final decision-makers, have full responsibility for medical outcomes, while AI systems themselves are not held accountable. This issue was also raised in our study. Specifically, if the AI system triggers an emergency alert (e.g., indicating a patient might have suicidal ideation), clinicians face multiple constraints in responding. These constraints might include a lack of resources or time to address the alert promptly, or situations where clinicians fail to review the AI system’s information in a timely manner. Any subsequent negative outcome, such as a life-threatening incident, raises the question of who should be held accountable.
We believe that in a high-risk and high-uncertainty medical scenario where the patients are young, it’s a key challenge to ensure the factuality of emergency alerts and find a balance between handling alerts and clinicians’ responsibilities throughout the development lifecycle of an AI system. Addressing this challenge requires collaboration among a broader set of stakeholders. For instance, involving the youth patient’s family in handling emergencies could play a crucial role. If the AI system detects an emergency, such as suicidal ideation, the system should notify not only the clinician but also the patient’s family immediately because family members could take appropriate action in cases where the clinician cannot respond promptly. For example, family members could calm the patient, take them to a healthcare facility, or contact mental health professionals for intervention. Additionally, clinicians who receive such alerts could collaborate with the family by providing timely professional guidance. Such notification and collaboration mechanisms would not only enable families to gain a more comprehensive understanding of the patient’s condition but also reduce risks arising from delayed clinical responses. Engaging parents in the emergency care-providing process ensures patient safety while alleviating some of the pressure on clinicians who might be constrained by limited time or resources.
To address ambiguity in responsibility for medical decisions, collaborative efforts involving hospital administrators, legal experts, and researchers are needed to develop clear guidelines on accountability in AI-supported clinical scenarios. Stakeholders should clearly define roles and accountability in emergencies triggered by AI systems to address the potential liability concerns highlighted in existing studies (Goodman and Flaxman, 2017). An addition benefits would be helping clinicians feel more secure when using AI systems in their clinical workflow.
Future research should explore mechanisms to promote effective collaboration between clinicians, parents, AI system developers, HCI designers, and other stakeholders. The multi-stakeholder collaboration will be essential for designing accountable and trustworthy AI-RPM systems that can handle high-risk scenarios while ensuring patient safety and clinician confidence.
7.3. AI as a Collaborative Decision-Making Partner For Patient Care Outside the Clinic
Our study highlights that concussion clinicians found certain AI-supported metrics—such as physical activity, sleep patterns, and nutritional intake—useful. The metrics about physical activities serve as critical indicators that clinicians interpret in relation to potential mental health sequelae. For instance, disruptions in sleep or drastic reductions in physical activity may signal the need for closer monitoring or further evaluation of youth concussion patients. However, some clinicians expressed reluctance to trust AI-generated probabilistic scores unless they were accompanied by high levels of accuracy and interpretability. Instead, they prefer using the AI system as a data collection and consolidation tool but leave clinical decisions to human experts.
This finding emphasizes the importance of designing AI-powered systems that do not attempt to replace clinical judgment but rather complement it by providing structured and reliable data that clinicians need to support their decision-making. As identified in previous CSCW studies on human-AI collaboration in high-stakes environments, trust and interpretability are key components influencing whether end-users effectively adopt AI systems in their workflows (Liao et al., 2020; Amershi et al., 2019). Our study explores how these principles apply in the specific context of youth concussion care, where clinicians balance multiple sources of uncertainty, patient variability, and time-sensitive decisions with respect to patients’ mental health conditions.
Our findings resonate with ongoing discussions in the CSCW community regarding the design of collaborative systems that support human decision-making while preserving user agency. Existing works emphasize that human-centered AI systems should aim to act as “intelligent assistants” rather than autonomous decision-makers (Yang et al., 2019; Zhang et al., 2020). In particular, enabling users to maintain control over critical decisions is critical in clinical settings where over-reliance on AI-generated recommendations could lead to adverse outcomes if predictions are incorrect or poorly contextualized. Our work contributes new insights on how AI-RPM systems can be effectively integrated into clinical workflows without introducing additional cognitive burdens or reducing clinician autonomy. Specifically, our study extends the understanding of human-AI collaboration by demonstrating how AI can facilitate patient care outside the clinic through continuous monitoring of non-clinical patient data to bridge the long-existing patient care gaps between clinical visits. While prior HCI and CSCW research has explored AI’s role in inpatient or acute care settings (Liao et al., 2020), we shift the focus to outpatient care and remote monitoring as an emerging important healthcare area.
Our study highlights the need for AI system design approaches that prioritize transparency, explainability, and collaboration across multiple stakeholders. Our findings suggest that AI-generated insights should be designed as part of an interactive process, where clinicians can not only access key metrics but also contextualize them with their broader understanding of a patient’s condition. For instance, future work could investigate interactive dashboards that allow clinicians to adjust or refine AI predictions based on their observations, creating a feedback loop between human expertise and machine learning outputs. The human-AI collaboration paradigm should not be a static and standardized format, instead, such collaboration should be seen as dynamic and iterative (Muller and Weisz, 2022; Fragiadakis et al., 2024).
For the information presentation strategies to support clinicians’ professional workflow, our study highlights the potential for collaborative data presentation strategies that tailor information to the specific needs of different stakeholders, including clinicians, patients, and caregivers. As suggested by prior HCI and CSCW studies on personalized information delivery (Amershi et al., 2019), layering information hierarchically and providing context-specific views can reduce information overload while ensuring that users have access to relevant data when needed. In the context of AI-RPM systems, we propose to display summary statistics for routine monitoring while offering detailed visualizations or explanations for cases requiring further investigation.
Our study not only provides practical recommendations for designing AI-RPM systems but also contributes to broader CSCW discussions on human-AI collaboration in high-stake and high-risk domains. Future research could explore collaborative design methodologies that involve clinicians, patients, and caregivers throughout the development process, ensuring that AI systems meet the diverse needs of all stakeholders.
Moreover, addressing trust concerns and ensuring interpretability remain key areas for future exploration (Barnes and Hutson, 2024; Mendel et al., 2024; Prahl and Van Swol, 2017; Dietvorst et al., 2015; Wang et al., 2019). The CSCW community could investigate how interactive system features, such as real-time feedback or explanatory interfaces, influence user trust and decision-making in long-term monitoring scenarios. Additionally, as AI-RPM systems become more prevalent, interdisciplinary collaborations involving legal, ethical, and technical perspectives will be necessary to establish guidelines for responsible deployment and use.

### v2

@@ 6.2. The Sleep and Physical Activities Module
In addition, clinicians requested the monitoring of additional metrics, such as heart rate during sleep and nap duration. To incorporate these data points without increasing cognitive load, we organized the data into three tabs: Sleep Data, Physical Activities, and Recommendation Compliance. The Sleep Data tab is presented by default, which allows clinicians to immediately access important metrics such as heart rate, nap duration, and sleep and wake-up times. They can choose to view one or multiple metrics, depending on their specific needs. Different data types are displayed in distinct colors to help clinicians visually distinguish between them more easily. If clinicians want to delve deeper into a patient’s sleep patterns, they can hover over a data point in the chart, then a more detailed value for that day will be displayed.
6.3. The Self-Report Symptoms Module
To maintain consistency with the Sleep and Physical Activities module, we added question marks indicating the data source and provided time selection options in the Self-Report Symptoms module (Fig. 1 D). Moreover, we identified emergent situations (i.e., suicidality) as both critical and challenging for clinicians to monitor, given resource constraints and the risk of false alarms. To address this, clinicians emphasized the importance of including parents in managing emergencies. Thus, we introduced two sections within this module. On the left side, we included the CA-parent conversation history, allowing clinicians to evaluate the patient’s health conditions from family perspectives. This conversation history can be expanded by clicking the icon next to the title ”Conversational History”. In addition, we added two call buttons at the bottom of the module, which enables clinicians to contact either the patient or a parent to manage emergency.
Besides, clinicians confirmed the value of different categories (i.e., depression, anxiety) in this module but indicated the need for more information. In response, we retained the original categories and introduced new ones—such as nutrition level—based on clinician feedback. Moreover,clinicians mentioned about the importance of severity for their treatment plan. To meet this need, we placed color-coded dots next to each symptom category (red for severe symptoms, orange for moderate, and green for none). To the right of each category, we provided a symptom duration indicator to show how long the reported issue has persisted, which help clinicians form a better understanding of patients health conditions. Furthermore, clinicians can click AI Summary to view a detailed breakdown of the category, providing additional context to support their assessment.
6.4. The Questionnaires Result Module
@@
In this module (Fig. 1 E), the title includes an option to view data source information, which comes from patient-administered questionnaires at home. These questionnaires include GAD-7 for anxiety, PHQ-9 for depression, and PCSS for concussion severity. The middle section of the chart presents the scores for each questionnaire, along with corresponding explanations to help clinicians interpret the results. The module also displays a comparison with previous scores, providing clinicians with a sense of the patient’s recovery trend. Additionally, clinicians can click ”Show Details” to view a breakdown of individual section scores within each questionnaire for a more in-depth assessment.
6.5. The AI Risk Score Prediction Module
The final module (Fig. 1 F) presents the AI-generated prediction of a patient’s risk of developing mental health sequelae. Given that concussion clinicians primarily manage patient care during the first four weeks following a concussion, we adjusted the default timeframe for risk predictions from one year to four weeks, while still providing other timeframe options for selection. Below the imeframe options, the visualization of the risk score chart remains unchanged from the preliminary design, given that clinicians reported it to be easy to interpret. However, clinicians expressed an interest in knowing which modules or features influence the AI-generated risk score. To highlight the relationship between the AI risk score and other modules, we introduce Features Contributing to the AI Risk Score section at the bottom of the module. The section includes contributing features and its corresponding module, along with a contribution rate. Moreover, we provide visual feedback on each feature’s contribution to the AI risk score. Both the color of the dots next to the feature and the font color of its associated percentage value dynamically change based on the degree of the contribution. In this way, clinicians can identify the important metrics quickly.
@@ 7. Discussion
In the following sections, we first discuss the use of AI-RPM technologies in the remote monitoring of patient mental health symptoms. Then, we talked about collaborative design in AI-RPM systems for enhancing usability and effectiveness. Finally, we conclude with collaborative emergency management with AI-RPM systems. Our work contributes to CSCW and HCI research at the intersection of health, RPM, and human-AI decision-making.
7.1. Unraveling Patients’ Intertwined Mental Health and Concussion Symptoms
@@
In clinical scenarios, it is not uncommon to have multiple illnesses intertwined with each other. For instance,  Wu et al. (2024) discovered that cancer oncologists need to collaborate with cardiologists to adjust the cancer treatment plan and be constantly aware of patients’ cardiotoxicity. Nevertheless, the concussion-mental health scenario and the cancer-cardiotoxicity scenario are fundamentally different. Concussion clinicians, according to our study, do not collaborate with mental health specialists, and neither do they diagnose mental health disorders. In most clinical practices, concussion clinicians assess whether a patient’s mental health symptoms exceed a certain severity threshold during in-person visits. Based on this assessment, they adjust the treatment plan accordingly.
We believe that AI-RPM technologies have the potential to give clinicians insights into the severity of patients’ mental health symptoms. First, LLM-based CAs could provide insight into patients’ mental health symptoms, such as the level of anxiety. Moreover, AI-based predictive models could help assess the likelihood of developing mental health sequelae, which can offer valuable and actionable insights into patients’ mental health conditions. As suggested by Bennett et al. (2012); Zhang et al. (2024), utilizing AI predictive models to provide actionable insights can increase the usefulness of AI in clinicians’ workflow. Future research could explore the design of AI-RPM technologies for collecting and distinguishing youth patients’ overlapping symptoms.
7.2. Augmenting Objective Data With Subjective Experience
Traditional RPM technologies, such as wearables, have already been used to remotely monitor concussion patients’ objective data (Yang et al., 2020). However, collecting objective data alone is not enough in our scenarios. As we discussed in the last section, it is critical to monitor the severity of mental health symptoms, such as anxiety and suicidal ideation. However, these symptoms are more subjective experiences that can hardly be captured by traditional RPM technologies. To address this gap, we propose using LLM-based CAs to gather patients’ anxiety levels and other mental health conditions at home. LLM-based CAs could automatically detect key psychological risk factors such as suicidal thoughts from patients’ conversations and generate a summary reportfor the day, which has been demoed in various prior research works in other scenarios (Bartle et al., 2023, 2022; Wang et al., 2023a; Simpson et al., 2020; Ma et al., 2024; Yang et al., 2023).
We believe that subjective information collected from LLM-powered CAs serves not only as supplementary information but also as contextual information for objective data (i.e., heart rate). Prior HCI and CSCW research has highlighted the principle that technology design should incorporate context into data (Dourish, 2004; Yoo et al., 2024). By combining subjective information and objective data in AI-RPM systems, clinicians could use the system to assess patients’ situations more accurately and avoid unnecessary concern over what might otherwise appear to be worrisome in objective metrics alone.
7.3. Collaborative Design in AI-RPM Systems for Enhanced Usability and Effectiveness
When presenting objective and subjective data in an AI-RPM system, it is important to prioritize the data presentation to the usability of the system. Concussion clinicians are already overwhelmed by existing EHR systems, and introducing an AI-RPM system with additional data could further increase their workload and strain limited healthcare resources. Nevertheless, clinicians expressed a willingness to briefly review AI-RPM data before each visit to better prepare for patient consultations. To provide information without increasing workload, designers need to collaborate with clinicians to prioritize data based on the importance of AI-RPM systems. Key data should be upfront while keeping secondary information hidden but accessible. Our work offers a refined system design that presents a clear information hierarchy in a visually intuitive manner (Fig. 1), serving as a reference for future development.
Moreover, domain knowledge is essential for defining thresholds of mental health symptom severity in AI-RPM system design. Without clear definitions of severity, AI-RPM systems may either underreport or overemphasize symptoms, leading to delayed interventions or unnecessary concerns. Thus, researcher and designers should closely collaborate with clinicians to establish clear thresholds for the severity of mental health symptoms so that clinicians get alert by the system only when necessary. Such a collaborative approach among clinicians, researchers, and designers could enhance the usability and effectiveness of clinician-facing AI-RPM systems. In the future, researchers and designers should focus on the key thresholds that influence clinician decisions.
7.4. Collaborative Emergency Management With AI-RPM Systems
Beyond the design of AI-RPM systems, it is also important to consider how the systems support clinicians during emergency situations in real-world practice.
In our study, concussion clinicians view AI as an important collaborative partner to support their decisions, which aligns with precious research in HCI and CSCW communities (Zhang et al., 2024; Hao et al., 2024; Yang et al., 2019; Zhang et al., 2020), However, they expressed two main concerns in emergencies. The first concern is the potential false alert of ”suicidal ideation” in the system. Youth patients may often experience mood swings and express emotions with exaggeration, which could lead to false alarms. As a result, a false alert could waste clinicians’ valuable time and critical medical resources that could be allocated to other patients. Moreover, if false alarms occur frequently, clinicians could lose trust in the system, which eventually leads to the abandonment of such a system (Liao et al., 2020; Amershi et al., 2019). Another concern is the ambiguity of accountability that AI-RPM systems introduce. When an AI-RPM system triggers an emergency alert (e.g., suicidal ideation), clinicians may face time or resource constraints that delay their response. Such delays can lead to serious outcomes and raise questions about accountability, which may affect system adoption.
In high-risk, uncertain cases involving youth, it’s challenging to verify alerts and balance response with accountability. We believe that addressing these two challenges requires collaboration among a broader set of stakeholders and clarifying accountability (Goodman and Flaxman, 2017) before implementing the system. Previous HCI research has primarily focused on involving clinicians and patients in emergency management with AI systems (Wu et al., 2024; Hao et al., 2024). We propose involving the youth patient’s family in the decision-making process alongside clinicians to manage emergencies. If suicidal ideation is detected, both the clinician and the patient’s family should be notified. Since family members are often nearby, they can respond quickly, while clinicians provide timely professional guidance. Such collaboration helps mitigate risks from delayed responses while ensuring shared responsibility in emergencies. Future research could explore parent-facing systems that integrate with clinicians’ systems to enhance remote youth patient care.